OpenStudy (anonymous):

Find the derivative of Y with respect to x: Y=10e^(-X/5). Can anybody help me with this prblem?

myininaya (myininaya):

okay so if u is a function of x we have the derivative of e^u=u'*e^u

myininaya (myininaya):

so y'=10*(-x/5)'e^(-x/5)

myininaya (myininaya):

=10*(-1/5)*e^(-x/5)

myininaya (myininaya):

=-2e^(-x/5)